OpenStudy (anonymous):

log(3x) - log(5) = 1

OpenStudy (anonymous):

are you solving for x? if im not mistaken this rule applies to your problem logb(m/n) = logb(m) – logb(n) so log(3x) - log(5) = 1 is log( 3x/5) = 1 and 3x/5=10^1=10

OpenStudy (anonymous):

yes Im solving for x

OpenStudy (anonymous):

3x/5=10 and now i think you got it from here right?

OpenStudy (anonymous):

how did you get the 10?

OpenStudy (anonymous):

oh wait I see

OpenStudy (anonymous):

THE ANSWER IS 16.66666667

OpenStudy (anonymous):

i got that too! thanks!!

OpenStudy (anonymous):

i have one more question

OpenStudy (anonymous):

2log(x) - log(3) = 1

OpenStudy (anonymous):



OpenStudy (anonymous):